# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v2

@@ 2.1.4 Automatic solution interpolation
SymBoltz integrates the background and several perturbation ODEs in conformal time $\tau$ for several wavenumbers $k$. All the results are stored in a special solution object. This solution object can be queried for any variable or symbolic expression
@@
This looks up the background solution if $S$ is a background variable, or the perturbation solutions if it is perturbative. It interpolates from the times stored by the ODE solver to the requested $\tau$ using the solver’s dense interpolation. If $S$ is perturbative, it also interpolates from the solved perturbation $k$-modes to the requested $k$. If $S$ is an unknown, it is returned directly from the ODE solution. If $S$ is observed, it is instead recomputed automatically from the equation that defines it as a function of the unknowns.
The result is that the user can access any variable in the model without having to do this manually. The solution interpolation is also incorporated into plot recipes for visualization of any variable as a function of $\tau$ and/or $k$.
@@ 2.1.5 Automatic Jacobian generation and sparsity detection
Just like SymBoltz generates code for $\boldsymbol{f}$ in the ODE (3), it uses the same symbolic equations to generate its Jacobian $\boldsymbol{J}$ with entries
@@
Jacobians are crucial for solving stiff ODEs with implicit solvers. Manually coding them is tiresome, error-prone, and must be repeated for new models. Numerical evaluation with finite differences is both approximate and slow. It can also be computed with automatic differentiation, but we find this slower and overkill compared to the analytical $\boldsymbol{J}$. Analytical Jacobians are as fast and stable as possible, and automatic generation lets the solver benefit from it while the modeler focuses on the high-level equations.
SymBoltz stores the Jacobian in sparse form to boost performance when it has many zeros. From the analytical $\boldsymbol{J}$, SymBoltz precomputes its exact sparsity pattern (where $J_{ij}=0$). This is hard to specify manually, and numerical solvers cannot distinguish false local zeros (for some input) from true global zeros (for all inputs). Without approximations, SymBoltz then reuses the same sparsity structure throughout the integration, computes the nonzero $J_{ij}$ analytically and stores them directly in the sparse $\boldsymbol{J}$ with no extra cost from looking up sparse matrix indices.
Easy, accurate, fast and sparse Jacobians are crucial for SymBoltz to remain performant without approximations! It is a major advantage with the symbolic approach, as in PyCosmo.
@@ 2.2 Approximation-freeness
SymBoltz treats stiffness in the Einstein-Boltzmann equations with modern implicit ODE solvers that integrate the full equations at all times. It is therefore free of approximation schemes, such as the TCA, UFA, RSA, NCDMFA and Saha approximation. This is friendlier to the modeler, who now simply has to provide a single set of equations instead of deriving approximations, validating them and dealing with related chores. This pairs well with a high-level equation-oriented symbolic interface.
@@
Implicit solvers take more expensive steps than explicit methods. At every step, they solve a generally nonlinear system of equations for the next unknowns $\boldsymbol{u}$. Using Newton’s method, it iteratively solves linear systems $\boldsymbol{W}\boldsymbol{u}=\boldsymbol{b}$ by LU-factorizing $\boldsymbol{W}=\boldsymbol{I}-\gamma h\boldsymbol{J}$, where $\gamma$ is a constant, $h$ is the time step and $\boldsymbol{J}$ is the ODE Jacobian (5). [^10] For example, the implicit Euler method $\boldsymbol{u}_{n+1}=\boldsymbol{u}_{n}+h\boldsymbol{f}(\tau_{n+1},\boldsymbol{u}_{n+1})$ solves $\boldsymbol{F}(\boldsymbol{u}_{n+1})=\boldsymbol{u}_{n+1}-h\boldsymbol{f}(\tau_{n+1},\boldsymbol{u}_{n+1})-\boldsymbol{u}_{n}=\boldsymbol{0}$ at every step with Newton’s method using its Jacobian $\boldsymbol{W}=\boldsymbol{\nabla}\boldsymbol{F}=\boldsymbol{I}-h\boldsymbol{J}(\tau_{n+1},\boldsymbol{u}_{n+1})$. LU-factorizing a dense $n\times n$ matrix from an $n$-dimensional ODE costs $O(n^{3})$ operations and bottlenecks larger ODEs (e.g. larger $l_{\text{max}}$). In return, implicit methods use more information about the ODE to take stable and long steps.
@@
To be fast, implicit solvers need several optimizations. They must compute not only $\boldsymbol{f}$, but also $\boldsymbol{J}$ efficiently. A common trick is to reuse an LU-factorized $\boldsymbol{W}$ over several steps, even if it really changes. Newton’s method only needs an approximate $\boldsymbol{W}$ to find the root, and $\boldsymbol{W}$ is recomputed only if convergence is slow due to an outdated $\boldsymbol{J}$ or a new step size $h$. This trades fewer LU-factorizations and $\boldsymbol{J}$ evaluations for more linear solve back-substitutions, which only cost $O(n^{2})$ operations. The next trick is to speed up the linear algebra with sparse matrices if $\boldsymbol{W}$ has many zeros, as for the perturbations. SymBoltz handles this by generating the analytical and sparse Jacobian (see section 2.1.5).
@@
SymBoltz generates code for all implicit solvers in OrdinaryDiffEq.jl [^11] https://github.com/SciML/OrdinaryDiffEq.jl/ and dense and sparse matrix methods in LinearSolve.jl [^12] https://github.com/SciML/LinearSolve.jl. We can test different solvers, which is useful as performance and stability of implicit methods and linear algebra operations are problem-dependent. Using $O(100)$ perturbation equations with more than $95\%$ sparsity, we find RFLUFactorization fastest for dense matrices, but KLUFactorization several times faster when exploiting sparsity. We tested these ODE solvers:
@@
  • Backward Differentiation Formula (BDF) methods like FBDF and QNDF (Shampine and Reichelt, 1997) are multi-step methods that use several past points to solve for the next step. They do not have Runge-Kutta ?stages? and solve just one implicit equation per step, so they scale well to large ODEs and can reuse Jacobians aggressively. They have variable order of 1-5, but are ?L-stable? only up to 2nd order. This makes them take small steps in stiff regimes without utilizing their high order. We find them to be slowest. The solvers ndf15 in CLASS and BDF2 in PyCosmo belong to this family. Nadkarni-Ghosh and Refregier (2017) also discuss BDF methods in the context of the Einstein-Boltzmann equations.
@@
  • Explicit Singly-Diagonal Implicit Runge-Kutta (ESDIRK) methods (e.g. Jørgensen et al., 2018) are a class of implicit Runge-Kutta methods designed to preserve most stability properties of Fully Implicit (FIRK) methods at a cheaper computational cost. An $s$-stage FIRK method has excellent stability, but a dense Butcher tableau $a_{ij}$ and must solve a system of $n\times s$ equations at every step. D in ESDIRK means that $a_{ij}$ has only lower triangular and diagonal nonzeros, decoupling the system into $s$ systems of $n$ equations that are faster to solve sequentially. S means that all diagonal nonzeros $a_{ii}=\gamma$ are equal, so all stages share the same $\boldsymbol{W}$-matrix and only one must be factorized and can be reused as in BDF methods. E means that $a_{11}=0$, so the first stage is explicit and cheap. Unlike BDF methods, they can remain L-stable without losing order to stiffness. We find the 4th order KenCarp4 method (Kennedy and Carpenter, 2003) to perform well with much fewer time steps. Bolt.jl also defaults to this solver. DISCO-EB uses Kvaerno5, but we find it slower.
@@
  • Rosenbrock methods: Rosenbrock methods (e.g. Lang, 2020) linearize $\boldsymbol{f}\approx\boldsymbol{J}\boldsymbol{u}$ and effectively replace Newton’s method with one linear solve. Their major upside is being free from convergence on a nonlinear problem. This suits the perturbation ODEs, as they are already linear. The downside is that they can need a new $\boldsymbol{J}$ at every step, which is slow to evaluate numerically in many problems. But generating the Jacobian analytically eliminates this issue. A subclass called ?Rosenbrock-W? methods can reuse Jacobians, but this strategy is not yet used by OrdinaryDiffEq.jl. We find the 5th order L-stable Rodas5P method (Steinebach, 2023) to be both most efficient and stable on both the background and perturbations, and have not found it used by other Boltzmann codes.
Performance of these methods is compared in section 4.
@@
Traditional codes use approximations to reduce stiffness and increase performance by evolving fewer ODE states when possible. Earlier works find marginal speedups from the TCA and UFA compared to using implicit solvers, but the RSA can significantly speed up models with high $l_{\text{max}}$ (Lesgourgues and Tram, 2011; Moser et al., 2022; Hahn et al., 2024). However, approximation-based codes must account for the structure of the equations to change, adding overhead to ODE solvers and restructuring sparse Jacobians. Approximation-free solvers can challenge this by optimizing around constant structure. Numerical codes approximate $\boldsymbol{J}$ with finite differences using $O(n)$ evaluations of $\boldsymbol{f}$. This is wasteful, as $\boldsymbol{f}=\boldsymbol{J}\boldsymbol{u}$ for linear perturbations and computing the nonzeros of $\boldsymbol{J}$ requires fewer operations than $\boldsymbol{f}$! Symbolic codes can take advantage of such properties. Approximations also complicate the code and put more load on modelers to derive, implement and validate them, and repeat the process for modifications that reintroduce stiffness or invalidate approximations.
@@ 2.3 Differentiability
SymBoltz is compatible with automatic differentiation. It can compute derivatives of any output quantity (e.g. ODE variables, $P(k)$ or $C_{l}$) with respect to any input parameters (e.g. $h$ or $\Omega_{c0}$). For example, it can be used to learn the output’s sensitivity to the input, perform Fisher forecasts or compute the Jacobian in the shooting method when parametrizing models with integrated variables. It can also compute the ODE Jacobian, but we do this symbolically to get it analytically with its sparsity pattern.
Currently, SymBoltz and Bolt.jl both work only with forward-mode dual numbers through ForwardDiff.jl [^13] https://github.com/JuliaDiff/ForwardDiff.jl. This is appropriate for applications with more outputs than inputs, like differentiating $O(100)$ values of $P(k)$ or $C_{l}$ as functions of $O(10)$ parameters. Reverse-mode is ideal with fewer outputs and attractive for MCMCs or training emulators where output is compressed to a scalar loss. Fast reverse-mode gradients could accelerate sampling in large parameter spaces from upcoming surveys and challenge use of emulators. Notably, DISCO-EB supports reverse-mode, but has not yet demonstrated it for this purpose. However, reverse-mode in SymBoltz is left for future work.
@@ 3 Examples
The best way to illustrate the features in section 2 is perhaps through some examples. The code is as of SymBoltz version 0.14.1 and available in a notebook in the project repository.
@@ 3.1 Basic usage workflow
Figure 2: Included plot recipes in SymBoltz make it trivial to visualize any symbolic variable or expression thereof from a solution of the Einstein-Boltzmann equations. This plot was made with one short line of code per subplot. Wavenumbers $k$ are in units of $H_{0}/c$.
@@
Most usage follows a model–problem–solution workflow: {codebox}
@@
using SymBoltzM = ΛCDM(lmax = 16)p = Dict( M.γ.T₀ => 2.7, M.b.Ω₀ => 0.05, M.b.YHe => 0.25, M.ν.Neff => 3.0, M.c.Ω₀ => 0.27, M.h.m_eV => 0.02, M.I.ln_As1e10 => 3.0, M.I.ns => 0.96, M.g.h => 0.7)prob = CosmologyProblem(M, p; jac=true, sparse=true)ks = [4, 40, 400, 4000] # k / (H₀/c)sol = solve(prob, ks)The model–problem–solution split achieves three distinct goals.
@@
First, a purely symbolic representation M of the $\mathrm{\Lambda CDM}$ model is created. This is a standalone object designed to be interactively inspected and modified. It contains every variable, parameter and equation of the cosmological model structured as one submodel per physical component: the metric $g$, the gravitational theory $G$, photons $\gamma$, massless neutrinos $\nu$, massive neutrinos $h$, cold dark matter $c$, baryons $b$, the cosmological constant $\Lambda$ and inflation $I$. For example, equations(M) shows all equations; equations(M.G) only the gravitational ones; M.g.a refers to the scale factor variable $a(\tau)$ that ?belongs? to the metric $g$; M.b.Ω₀ is the baryon density parameter $\Omega_{b0}$; and parameters(M) lists parameters that can be set. This displays with -compatibility in notebooks and transparently shows the equations being solved.
@@
Second, the symbolic model is compiled to a numerical problem prob with parameters p. This is an expensive step where the processing in section 2.1 happens: equations are checked for consistency and split into background/perturbation stages; observed/unknown variables are distinguished; fast ODE code is generated; background unknowns in the perturbations are replaced by splines; the Jacobian matrix is generated in numerical/analytical and dense/sparse form; and initial values are computed. The compilation is customized with optional arguments like jac and sparse. This is a separate step because it performs final transformations on the model M once the user has finished modifying and committed to it.
@@
Third, solve(prob, ks) solves the background and perturbations for the wavenumbers ks. Omitting ks solves only the background. The resulting solution object sol provides convenient access to all model variables. Internally, it stores values of all ODE unknowns at the time steps taken by the solvers and for every requested wavenumber. But it can be queried with any time, wavenumber and symbolic expression, and will automatically compute it from the unknowns and interpolate between stored times and wavenumbers, as explained in section 2.1.4. For example, calling sol(1.0, 2.0, g.Φ+g.Ψ) computes $\Phi+\Psi$ at $k=1\,H_{0}/c$ and $\tau=2\,H_{0}^{-1}$ by expanding it in terms of unknowns using the equations in section A.2. This interface was used to plot the evolution of several variables in fig. 2 with little code.
@@
Most Boltzmann solvers save only a fixed set of variables, such as only the unknowns. Recovering observed variables requires manual effort. This is cumbersome, opens for user error and adds friction in the modeling process. SymBoltz is designed to provide easy access to all variables defined by the model.

### v3

@@ 2.1.4 Automatic solution interpolation
SymBoltz integrates the background and perturbation ODEs in conformal time $\tau$ and for several wavenumbers $k$. All the results are stored in a dedicated solution object. This object can be queried for any variable or symbolic expression
@@
This looks up the background solution if $S$ is a background variable, or the perturbation solutions if it is perturbative. It interpolates from the times stored by the ODE solver to the requested $\tau$ using the solver’s dense interpolation scheme. If $S$ is perturbative, it also interpolates from the solved perturbation $k$-modes to the requested $k$. If $S$ is an unknown, it is returned directly from the ODE solution. If $S$ is observed, it is instead recomputed automatically from its expansion in terms of unknowns.
The result is that the user can access any variable in the model without having to do any of this manually. The solution interpolation is also incorporated into plotting recipes that easily visualize any variable as a function of $\tau$ and $k$.
@@ 2.1.5 Automatic Jacobian generation and sparsity detection
Just as SymBoltz generates code for $\boldsymbol{f}$ in the ODE (3), it uses the same symbolic equations to generate its Jacobian $\boldsymbol{J}$ with entries
@@
Jacobians are crucial for solving stiff ODEs with implicit solvers. Manually coding them is tiresome, error-prone, and must be repeated for new models. Numerical evaluation with finite differences is approximate and slow. Automatic differentiation can compute $\boldsymbol{J}$ from $\boldsymbol{f}$, but makes it harder to exploit the sparsity of $\boldsymbol{J}$. Analytical Jacobians are as fast and stable as possible, and SymBoltz generates them automatically for the solver without distracting the modeler from focusing on the physical equations.
SymBoltz stores the Jacobian in sparse form to boost performance when it has many zeros. From the analytical $\boldsymbol{J}$, SymBoltz precomputes its exact sparsity pattern (where $J_{ij}=0$). This is hard to specify manually, and numerical solvers cannot distinguish false local zeros (for some inputs) from true global zeros (for all inputs). Without approximations, SymBoltz reuses the same sparsity structure throughout the integration, computes the nonzero $J_{ij}$ analytically, and stores them at a precomputed index in the sparse $\boldsymbol{J}$. This avoids converting $\boldsymbol{J}$ from dense to sparse form at runtime. Easy analytical and sparse Jacobians are crucial for SymBoltz to remain fast without approximations. This is a major advantage with the symbolic approach, as in PyCosmo.
@@ 2.2 Approximation-freeness
SymBoltz treats stiffness in the Einstein–Boltzmann equations with modern implicit ODE solvers that integrate the full equations at all times. It is therefore free of approximation schemes, such as the TCA, UFA, RSA, NCDMFA, and Saha approximation. This is friendlier to the modeler, who now has to provide only one set of equations, instead of deriving, implementing, and validating approximations. This approach is also well-suited to the high-level equation-oriented symbolic interface.
@@
Implicit solvers take more expensive steps than explicit methods. At every step, they solve a generally nonlinear system of equations for the next unknowns $\boldsymbol{u}$, often using Newton’s method. It iteratively solves linear systems $\boldsymbol{W}\boldsymbol{u}=\boldsymbol{b}$ by LU-factorizing $\boldsymbol{W}=\boldsymbol{I}-\gamma h\boldsymbol{J}$, where $\gamma$ is a constant, $h$ is the time step, and $\boldsymbol{J}$ is the ODE Jacobian (5). [^16] For example, the implicit Euler method $\boldsymbol{u}_{n+1}=\boldsymbol{u}_{n}+h\boldsymbol{f}(\tau_{n+1},\boldsymbol{u}_{n+1})$ solves $\boldsymbol{F}(\boldsymbol{u}_{n+1})=\boldsymbol{u}_{n+1}-h\boldsymbol{f}(\tau_{n+1},\boldsymbol{u}_{n+1})-\boldsymbol{u}_{n}=\boldsymbol{0}$ at every step with Newton’s method using its Jacobian $\boldsymbol{W}=\boldsymbol{\nabla}\boldsymbol{F}=\boldsymbol{I}-h\boldsymbol{J}(\tau_{n+1},\boldsymbol{u}_{n+1})$. LU-factorizing a dense $n\times n$ matrix from an $n$-dimensional ODE costs $O(n^{3})$ operations and bottlenecks larger ODEs (e.g., larger $l_{\text{max}}$). In return, implicit methods use more information to take long and stable steps.
@@
To be fast, implicit solvers need several optimizations. They must compute not only $\boldsymbol{f}$, but also $\boldsymbol{J}$ efficiently. A common trick is to reuse an LU-factorized $\boldsymbol{W}$ over several steps, even if it really changes. Newton’s method only requires an approximate $\boldsymbol{W}$ to find the root, and $\boldsymbol{W}$ is recomputed only if convergence is slow due to an outdated $\boldsymbol{J}$ or a changed step size $h$. This trades fewer LU-factorizations and $\boldsymbol{J}$ evaluations for more linear solve back-substitutions, which only cost $O(n^{2})$ operations. The next trick is to speed up the linear algebra with sparse matrices if $\boldsymbol{W}$ has many zeros, as for the perturbations. SymBoltz handles this by generating the analytical and sparse Jacobian (see Sect. 2.1.5).
@@
SymBoltz generates code for all implicit solvers in OrdinaryDiffEq, and dense and sparse matrix methods in LinearSolve. We can test different solvers, which is useful as performance and stability of implicit methods and linear algebra operations are problem-dependent. Using $O(100)$ perturbation equations with more than $95\%$ sparsity, we find RFLUFactorization in LinearSolve fastest for dense matrices, but KLUFactorization several times faster when exploiting sparsity. We tested these ODE solvers from OrdinaryDiffEq:
@@
  • Backward differentiation formula (BDF) methods, such as FBDF and QNDF (Shampine and Reichelt, 1997), are multi-step methods that use several past points to solve for the next. They do not have Runge–Kutta ?stages? and solve just one implicit equation per step, so they scale well to large ODEs and can reuse Jacobians heavily. They have variable order of 1-5, but are ?L-stable? only up to second order. This makes them take short steps in stiff regimes without utilizing their high orders. We find them to be slowest. The solvers ndf15 in CLASS and BDF2 in PyCosmo also belong to this family. Nadkarni-Ghosh and Refregier (2017) also discuss BDF methods in the context of the Einstein–Boltzmann equations.
@@
  • Explicit singly diagonal implicit Runge–Kutta (ESDIRK) methods (e.g., Jørgensen et al., 2018) are a class of implicit Runge–Kutta methods designed to preserve most stability properties of fully implicit (FIRK) methods at a cheaper computational cost. An $s$-stage FIRK method has excellent stability, but a dense Butcher tableau $a_{ij}$, and must solve a system of $n\times s$ equations at every step. D in ESDIRK means that $a_{ij}$ has only lower triangular and diagonal nonzeros, decoupling the system into $s$ systems of $n$ equations that are faster to solve sequentially. S means that all diagonal nonzeros $a_{ii}=\gamma$ are equal, so all stages share the same $\boldsymbol{W}$-matrix, and only one must be factorized and can be reused as in BDF methods. E means that $a_{11}=0$, so the first stage is explicit and cheap. Unlike BDF methods, they can remain L-stable without losing order to stiffness. We find the fourth order KenCarp4 method (Kennedy and Carpenter, 2003) to perform well with much fewer time steps. Bolt also defaults to it. DISCO-EB uses Kvaerno5, but we find it to be slower.
@@
  • Rosenbrock methods (e.g., Lang, 2020) linearize $\boldsymbol{f}\approx\boldsymbol{J}\boldsymbol{u}$ and effectively replace Newton’s method with one linear solve. Being free from nonlinear convergence issues is a major upside and suits the linear perturbation ODEs. The downside is that they can need a new $\boldsymbol{J}$ at every step, which is slow to evaluate in many problems. But generating $\boldsymbol{J}$ analytically eliminates this issue. The subclass of ?Rosenbrock-W? methods can reuse Jacobians, but this strategy is not yet used by OrdinaryDiffEq. We find the fifth order L-stable Rodas5P method (Steinebach, 2023) to be both most efficient and stable on both the background and perturbations, and have not found it used by other Boltzmann codes.
Section 4 compares the performance of these methods.
@@
Traditional codes use approximations to reduce stiffness and increase performance by evolving fewer ODE states when possible. Earlier works find marginal speedups from the TCA and UFA compared to using implicit solvers, but the RSA can significantly speed up models with high $l_{\text{max}}$ (Lesgourgues and Tram, 2011; Moser et al., 2022; Hahn et al., 2024). However, approximation-based codes must account for the structure of the equations to change, which adds overhead to ODE solvers and restructuring of sparse Jacobians. Approximation-free solvers can challenge this by optimizing around unchanged structure. Numerical codes approximate $\boldsymbol{J}$ with finite differences using $O(n)$ evaluations of $\boldsymbol{f}$. This is wasteful, as $\boldsymbol{f}=\boldsymbol{J}\boldsymbol{u}$ for linear perturbations, so computing the nonzeros of $\boldsymbol{J}$ should take fewer operations than $\boldsymbol{f}$! Symbolic codes can take advantage of such properties. Approximations also complicate the code and put more load on modelers to derive, implement, and validate them, and repeat the process as modifications reintroduce stiffness or invalidate approximations.
@@ 2.3 Differentiability
SymBoltz is compatible with automatic differentiation. It can compute derivatives of any output quantity (e.g., ODE variables, $P(k)$, or $C_{l}$) with respect to any input parameters (e.g., $h$ or $\Omega_{c0}$). For example, it can be used to learn the sensitivity of the output to the input, perform Fisher forecasts, or compute the Jacobian in the shooting method for boundary-value parameterized models. It can also compute the ODE Jacobian, but SymBoltz does this symbolically to get it analytically with its sparsity pattern.
Currently, SymBoltz and Bolt both work only with forward-mode dual numbers through ForwardDiff. This is appropriate for applications with more outputs than inputs, such as differentiating $O(100)$ values of $P(k)$ or $C_{l}$ as functions of $O(10)$ parameters. Reverse-mode is ideal with fewer outputs and attractive for MCMCs or training emulators where output is compressed to a scalar loss. Fast reverse-mode gradients could accelerate sampling in large parameter spaces from upcoming surveys and challenge the use of emulators. Notably, DISCO-EB supports reverse-mode, but has not yet demonstrated it for this purpose. However, reverse-mode in SymBoltz is left for future work.
@@ 3 Examples
The features in Sect. 2 are best illustrated through some examples. This code is as of SymBoltz version 1.0.0 and available in a notebook in the project repository.
@@ 3.1 Basic usage workflow
Figure 2: SymBoltz includes plotting recipes that make it easy to visualize any symbolic variable or expression thereof from a solution of the Einstein–Boltzmann equations. This plot was made with one short line of code per subplot. Wavenumbers $k$ are in units of $H_{0}/c$.
@@
Most usage follows a model-problem-solution workflow: {codebox}
@@
using SymBoltzM = ΛCDM(lmax = 16)p = Dict( M.γ.T₀ => 2.7, M.b.Ω₀ => 0.05, M.b.YHe => 0.25, M.ν.Neff => 3.0, M.c.Ω₀ => 0.27, M.h.m_eV => 0.02, M.I.ln_As1e10 => 3.0, M.I.ns => 0.96, M.g.h => 0.7)prob = CosmologyProblem(M, p; jac=true, sparse=true)ks = [4, 40, 400, 4000] # k / (H₀/c)sol = solve(prob, ks)The model-problem-solution split achieves three distinct goals.
@@
First, a symbolic representation M of the $\mathrm{\Lambda CDM}$ model is created. This is a standalone object designed to be interactively inspected and modified. It contains every variable, parameter, and equation of the model structured as one submodel per physical component: the metric $g$, gravitational theory $G$, photons $\gamma$, massless neutrinos $\nu$, massive neutrinos $h$, cold dark matter $c$, baryons $b$, the cosmological constant $\Lambda$, and inflation $I$. For example, equations(M) shows all its equations; equations(M.G) only the gravitational ones; M.g.a refers to the scale factor variable $a(\tau)$ that ?belongs? to the metric $g$; M.b.Ω₀ is the baryon density parameter $\Omega_{b0}$; and parameters(M) shows all parameters that can be set. Everything displays with -compatibility in notebooks and transparently shows the contents of the model.
@@
Second, the symbolic model is compiled to a numerical problem prob with parameters p. This expensive step does the processing in Sect. 2.1: it checks equations for consistency; splits them into background and perturbation stages; distinguishes observed and unknown variables; generates fast code for ODEs and Jacobians in numerical or analytical and dense or sparse form; splines background unknowns in the perturbations; and finds initial conditions. Optional keyword arguments customize the compilation and control precisely how the problem is solved. This is a separate step because it does final transformations on the model M once the user has finished modifying and committed to it.
@@
Third, solve(prob, ks) solves the background and perturbations for the wavenumbers ks. Omitting ks solves only the background. The resulting solution object sol provides convenient access to all model variables. Internally, it stores values of all ODE unknowns at the time steps taken by the solvers for every requested wavenumber. However, it can be queried with any time, wavenumber, and symbolic expression, and will automatically compute it from the unknowns and interpolate between stored times and wavenumbers, as explained in Sect. 2.1.4. For example, sol(g.Φ+g.Ψ, 1.0, 2.0) computes $\Phi+\Psi$ at $\tau=1\,H_{0}^{-1}$ and $k=2\,H_{0}/c$ by expanding $\Psi$ in terms of $\Phi$ and other unknowns. This interface is also used to plot the evolution of several variables in Fig. 2 with very compact code.
@@
Most Boltzmann solvers save only a fixed set of variables, such as only the unknowns. Recovering observed variables requires manual effort. This is cumbersome, open to user error, and adds friction in the modeling process. SymBoltz is designed to provide easy access to all variables defined by the model.